OpenStudy (anonymous):

kingvssj00 0 1/2 + 3/8 + 3/4 lesson: adding and subtracting unlike denomitators a minute ago

OpenStudy (anonymous):

convert into the same denominators: 4/8+3/8+6/8 =13/8 which is 1 5/8

OpenStudy (anonymous):

what

OpenStudy (anonymous):

you want to get a common denominator of 8 1/2 + 3/8 + 3/4 1/2 *(4/4) = 4/8 3/8 already has 8 as the denominator (3/4)*(2/2) = 6/8 so when you multiply by x/x its like multiplying by 1 3/4 + 4/8 + 6/8 = (3+4+6)/8 = ?